# NSR-Boost: A Neuro-Symbolic Residual Boosting Framework for Industrial Legacy Models

### v1

@@ A.3. Baseline Details
  • Res-XGBoost is a residual learning baseline constructed to verify the necessity of symbolic regression. This method directly uses gradient boosting trees to fit the residuals of the base model. In the experiments, we strictly limited the maximum tree depth to 3 to simulate the complexity of lightweight experts and introduced a Bayesian optimization strategy to automatically tune the learning rate and regularization terms.
@@ A.4. NSR-Boost Prompt Details
A.5. Additional Results
Table 3. Standard deviation of performance on various datasets.
@@ A.5. Additional Results
Experimental Stability Analysis. To evaluate the robustness of NSR-Boost against the stochasticity inherent in LLM generation and optimization processes, we report the performance standard deviations over five independent runs in Table 3. The results indicate that NSR-Boost maintains low variance levels, which are comparable to deterministic numerical baselines.
LLM Model Selection and Stability Analysis. As illustrated in Figure 11, our comprehensive analysis of failure modes across four distinct LLMs reveals a consistent pattern: the primary bottleneck in expert code construction lies not in syntactic correctness, but in semantic validity. The dominant failure category, Initial AUC Failure (the dark blue region), accounts for a substantial proportion of all rejected cases, ranging from approximately 27% to over 60% depending on the model. This phenomenon characterizes a form of “hallucination”, wherein the LLM-generated Python code is syntactically flawless and executable yet fails to capture the underlying residual patterns, resulting in no tangible improvement in validation AUC.

### v2

@@ A.3. Baseline Details
  • Res-XGBoost is a residual learning baseline constructed to verify the necessity of symbolic regression. This method directly uses gradient boosting trees to fit the residuals of the base model. In the experiments, we introduce a Bayesian optimization strategy to automatically adjust the learning rate, tree depth, and regularization term.
@@ A.4. NSR-Boost Prompt Details
Appendix B Additional Results
@@
Table 3. Standard deviation of performance on various datasets over five independent runs.
B.1. Experimental Stability Analysis
To evaluate the robustness of NSR-Boost against the stochasticity inherent in LLM generation and optimization processes, we report the performance standard deviations over five independent runs in Table 3. The results indicate that NSR-Boost maintains low variance levels, which are comparable to deterministic numerical baselines.
Figure 11. Visualization of the Symbolic Expert’s decision surface in Figure 3.
B.2. Interpretability
Figure 11 provides a geometric visualization of the symbolic expert logic presented in Figure 3 of the main text. This three dimensional surface bridges abstract mathematical formulas with their concrete behaviors within the decision space. The steep vertical truncation displayed within the red dashed box directly maps to the conditional statements found in the source code. This distinct visual feature provides compelling evidence that the model strictly enforces safety constraints by completely halting the correction mechanism for any applicants older than 34.3 years. Furthermore the smooth peak like elevation enclosed in the black dashed box offers an intuitive visual representation of the Gaussian interaction term defined within the algorithm. This geometric form confirms that the Symbolic Expert successfully identified a nonlinear opportunity window centered at 25.1 years wherein applicants possessing high repayment slack are recognized as high quality candidates and receive positive score corrections.
B.3. Performance on High-Value Customer Segment
To further verify the generalization capability of NSR-Boost across distinct business lines, we report supplementary deployment results from a third core industrial scenario targeting the High Value Segment of the B Card.
Table 4 details the monitoring data derived from processing approximately 1.41 million real world samples between February 2025 and July 2025. Consistent with the primary experimental setup, we maintained the Legacy Model in a frozen state. Experimental results demonstrate that NSR-Boost achieved sustained performance improvements within this scenario. Notably, NSR-Boost achieves a weighted average AUC improvement of 0.43% and a weighted average KS improvement of 0.92% during the strict OOT online deployment phase from May 2025 to July 2025.
Phase Month Legacy Model NSR-Boost
AUC KS AUC KS $\Delta$ AUC $\Delta$ KS
[…]
Note: The training data (Feb-25 to Apr-25) contains mixed validation samples, shuffled and split with an 80%/20% ratio.
Table 4. Performance evaluation of the High-Value Segment (B-Card) model. The online deployment processes approximately 1.41 million real user samples.
B.4. LLM Model Selection and Stability Analysis
As illustrated in Figure 12, we conduct a comprehensive analysis of the failure modes of four different LLMs, namely DeepSeek-R1 (DeepSeek-AI, 2025), Qwen3-30B-A3B-Thinking (Team, 2025b), NVIDIA-Nemotron-Nano-12B-V2 (NVIDIA, 2025), and Seed-OSS-36B-Instruct (Team, 2025a),reveals a consistent pattern: The primary bottleneck in expert code construction lies not in syntactic correctness, but in semantic validity. The dominant failure category, “Initial AUC Failure”, accounts for a substantial proportion of all rejected cases, ranging from approximately 27% to over 60% depending on the model. This phenomenon manifests as a form of hallucination where the LLM-generated Python code is syntactically flawless and executable yet fails to capture the underlying residual patterns, resulting in no substantive improvement in validation AUC.
@@ B.4. LLM Model Selection and Stability Analysis
B.5. Generalization Ability across LLMs.
Based on this rigorous verification mechanism, we further explore whether the effectiveness of NSR-Boost is bound to a specific LLM. We conduct a sensitivity analysis using the four aforementioned open-source LLMs, executing 20 independent runs for each model on the private financial dataset.
As shown in Table 5, all four backbones successfully enable the framework to outperform the initial baseline. It is worth noting that the Seed-OSS-36B model and DeepSeek-R1, which exhibit strong inference ability, achieve the top two results. The extremely small performance gap between different architectures proves that NSR-Boost is model-agnostic and robust, and can extract stable incremental gains regardless of the LLM. Based on these empirical results, especially considering the maximum improvement in performance, we ultimately choose Seed-OSS-36B.
LLM Model Average AUC Improvement
Legacy Model 0.688982 -
Nemotron-Nano-12B 0.689159 +0.000177
Qwen3-30B 0.689249 +0.000267
DeepSeek-R1 0.689302 +0.000320
Seed-OSS-36B (Ours) 0.689388 +0.000406
Table 5. Performance comparison of different LLMs on the private dataset.
B.6. In-depth Analysis of Residual Correction Benchmarks.
Method Paradigm AUC Gain Interp.
Legacy Model GBDT - High
Res-XGBoost Black-box -1.02e-3 Low
gplearn (20min) SR -1.78e-3 Moderate
gplearn (8.3h) SR $\pm$0.00e-3 Low
NSR-Boost Neuro-Symbolic +4.06e-4 High
Table 6. Performance and interpretability comparison of residual learning methods on the private OOT test set.
Table 6 presents a detailed comparison of residual correction performance across different paradigms on the OOT test set of the private dataset.
Purely data-driven baseline methods encountered significant generalization bottlenecks. As a representative black-box model, Res-XGBoost exhibited a negative gain of $1.02\text{e-}3$ despite undergoing extensive hyperparameter optimization via Optuna. This outcome indicates that given the robustness of the Legacy Model, the remaining residuals contain substantial random noise. Consequently, unconstrained strong estimators are prone to overfitting this noise, leading to a marked degradation in performance on the test set.
Traditional symbolic regression methods exposed their limitations regarding search efficiency in high-dimensional spaces. Under the standard setting of a 20-minute search, gplearn similarly succumbed to overfitting. More instructively, even after investing substantial computational resources for an intensive 8.3-hour search, the algorithm ultimately converged to a gain of $\pm 0.00$ with its final output being a trivial constant solution of -0.01. This phenomenon demonstrates that without the guidance of prior knowledge, genetic algorithms struggle to overcome the failure of search strategies within complex operator spaces caused by high-dimensional feature combinations.
In contrast, NSR-Boost stands out as the sole method in the experiment to achieve a positive gain of $4.06\text{e-}4$. By leveraging semantic understanding to effectively prune the search space, NSR-Boost enables the model to precisely locate feature interactions with clear business logic within extremely low signal-to-noise environments. This capability allows it to circumvent overfitting while simultaneously breaking through the performance ceiling inherent to purely data-driven methods.
Furthermore, regarding interpretability, although Res-XGBoost possesses strong fitting capabilities, its nature as a black-box ensemble composed of thousands of trees results in a total loss of physical interpretability. Meanwhile, Symbolic Regression in high search spaces tends to exhibit two extremes, yielding either unreadable long formulas due to data complexity or degenerating into trivial solutions to pursue generalization as shown in the table. Conversely, NSR-Boost strikes a balance between these extremes, achieving a unification of interpretability and performance.